# Rates of convergence in long time asymptotics of an alignment model with symmetry breaking

### v2

@@ 2.6. Log-Sobolev and Poincaré inequality
$\frac{d}{dt}\mathcal{F}[f(t,.)]=-\mathcal{I}[f(t,.)]\leq-\,D^{2}\,\mathcal{K}\,\mathcal{H}[f|G_{u_{f}}],$
where $\mathcal{K}=\mathcal{K}_{\epsilon}$ is the positive constant in (LSI) with $\epsilon=\sup_{t\geq 0}\mathrm{dist}(u_{f}(t),\mathcal{S})<+\infty$ by Lemma˜2.1. Since $t\mapsto\mathcal{F}[f(t,.)]$ is nonincreasing and bounded from below, it has a limit $\mathcal{F}_{\infty}\in\mathbb{R}$ as $t\rightarrow+\infty$. Integrating the previous inequality in time, we get a de Bruijn type estimate
$\int_{0}^{\infty}\mathcal{H}[f|G_{u_{f}}]dt\leq\frac{1}{\mathcal{K}D^{2}}(\mathcal{F}[f_{\mathrm{ini}}]-\mathcal{F}_{\infty}).$ (17)
With (16), (C-K) and (17), we obtain that $\gamma(t)=W(u_{f}(t))$ is integrable on $\mathbb{R}_{+}$, where $W$ is defined on $\mathbb{R}^{d}$ by $W(Y)=|Y-u_{G_{Y}}|^{\beta}$. Observe that $\gamma^{\prime}(t)=\nabla W(u_{f}(t))u_{f}^{\prime}(t)$ is bounded since $u_{f}$ and $u_{f}^{\prime}=\alpha u_{f}-\alpha\int_{\mathbb{R}^{d}}v|v|^{2}f\,dv$ are uniformly bounded in $t\geq 0$. Since $\gamma$ is integrable in time and uniformly continuous, we obtain $\lim_{t\rightarrow\infty}\gamma(t)=0$ thanks to Barbalat’s lemma. Using (7), we obtain :
$|u_{f}(t)-u_{G_{u_{f}}}(t)|=-\,\frac{{H}(|u_{f}|)}{{K}(|u_{f}|)}\frac{u_{f}}{|u_{f}|}.$
@@
$\lim_{t\rightarrow{\infty}}u_{f}(t)=0\,.$
If $D<D_{*}$, we recall that the zeros of H{H} on $\mathbb{R}_{+}$ are exactly $r(D)$ and 0 by Proposition˜1.1. If $\mathcal{F}[f_{\mathrm{ini}}]<\mathcal{F}[G_{0}]$, the case $\lim_{t\rightarrow{\infty}}u_{f}(t)=0$ is excluded, since in this case
$\mathcal{F}[f_{\mathrm{ini}}]\geq\lim_{t\rightarrow\infty}\mathcal{F}[f]\geq\inf_{u_{f}=0}\mathcal{F}[f]=\mathcal{F}[G_{0}]>\mathcal{F}[f_{\mathrm{ini}}].$
@@
$\lim_{t\rightarrow{\infty}}\mathrm{dist}(u_{f},\mathcal{S})=0\,.$
@@
$\mathcal{V}(u)=-\,D\log\left(\int_{\mathbb{R}^{d}}e^{-\frac{1}{D}\left(\frac{1}{2}|v-u|^{2}+\psi_{\alpha}(v)\right)}\,dv\right).$ (18)
In particular, $\mathcal{V}$ is smooth and radial and with a slight abuse of notation, we shall write $\mathcal{V}(u)=\mathcal{V}(|u|)$. The key point is to prove estimates of Polyak–Łojasiewicz type, which are based on the following Taylor expansion.
@@
  • If $D<D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{\prime\prime}(r(D))|<\epsilon$ for $i=1,2$ and $0<\delta<\mathrm{min}\big{\{}r(D)/2,\epsilon\big{\}}$ such that for all $u$ satisfying $\mathrm{dist}(u,\mathcal{S})<\delta$ we have,
    $\frac{\mu_{1}}{2}\mathrm{dist}(u,\mathcal{S})^{2}\leq\mathcal{V}(u)-\mathcal{V}_{*}\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(u)|^{2}.$ (19)
  • If $D>D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{\prime\prime}(0)|<\epsilon$ for $i=1,2$ and $0<\delta<\epsilon$ such that for all $|u|<\delta$ we have,
    $\frac{\mu_{1}}{2}|u|^{2}\leq\mathcal{V}(u)-\mathcal{V}(0)\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(u)|^{2}.$ (20)
  • If $D=D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{(4)}(0)|<\epsilon$ for $i=1,2$ and $0<\delta<\epsilon$ such that for all $|u|<\delta$ we have,
    $\frac{\mu_{1}}{24}|u|^{4}\leq\mathcal{V}(u)-\mathcal{V}(0)\leq\ \frac{6^{\frac{4}{3}}}{24\mu_{2}^{\frac{1}{3}}}|\nabla\mathcal{V}(u)|^{\frac{4}{3}}.$ (21)
@@
The proof follows from a Taylor expansion of $\mathcal{V}$ and $\mathcal{V}^{\prime}$ around $u\in\mathcal{S}$ if $D<D_{*}$ and around 0 if $D\geq D_{*}$. Since $\mathcal{S}$ and 0 are critical points of $\mathcal{V}$ (see Lemma˜3.2), one has to check that:
@@
It is shown in the proof of [11, Proposition 2.3] for the case $d=2$, and in the proof of [11, Proposition 2.2] for $d=1$, that
@@
We will prove that $H^{\prime}$ has a unique zero $u_{0}$ and $0<u_{0}<r(D)$. Since $D<D_{*}$, thanks to Proposition˜1.1, we know that $H(0)=0$ and $H(r(D))=0$ with $r(D)>0$. Using Rolle’s lemma, we conclude that $H^{\prime}$ has at least one zero on $(0,r(D))$.
@@ 3.2. Case $D<D_{*}$
Let $\epsilon>0$. By Lemma˜3.1 there exist positive constants $\mu_{1}$ and $\mu_{2}$ such that $|\mu_{i}-\mathcal{V}^{\prime\prime}(r(D))|<\epsilon$ and a neighborhood $\mathcal{N}=\mathcal{N}_{\epsilon}$ of $\mathcal{S}$ such that (19) holds for all $u\in\mathcal{N}$. By Lemma˜2.2, that is $\lim_{t\rightarrow\infty}\mathrm{dist}(u_{f}(t),\mathcal{S})=0$, there exists $t_{0}>0$ such that for all $t\geq t_{0}$ we have $u_{f}(t)\in\mathcal{N}$. Hence, we have,
$\mathcal{V}(u_{f}(t))-\mathcal{V}_{*}\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(u_{f}(t))|^{2}$ $=\frac{1}{2\mu_{2}}|u_{f}-u_{G_{u_{f}}}|^{2}$
@@
$\leq A_{q}(2\mathcal{H}[f(t,.)|G_{u_{f}(t)}])^{\frac{1}{p}}\text{ for all }t\geq t_{0},$ (24)
@@
$A_{q}=\frac{1}{2\mu_{2}}\sup_{t>t_{0}}{\left(\int_{\mathbb{R}^{d}}|v|^{q}|f+G_{u_{f}}|\,dv\right)^{\frac{2}{q}}}$
@@
$\mathcal{F}[f]-\mathcal{F}_{*}\leq\varphi(\mathcal{H}[f|G_{u_{f}}])\,\,\,\forall t\geq t_{0}$ (25)
@@
$\frac{d}{dt}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})=-\mathcal{I}[f(t,.)]\leq-D^{2}\mathcal{K}\mathcal{H}[f(t,.)|G_{u_{f}(t)}]\leq-D^{2}\mathcal{K}\varphi^{-1}\left(\mathcal{F}[f(t,.)]-\mathcal{F}_{*}\right),$ (26)
using (LSI) with constant $\mathcal{K}=\mathcal{K}_{\epsilon}$ since $\mathrm{dist}(u_{f}(t),\mathcal{S})<\epsilon$ for all $t\geq t_{0}$. Since $\mathcal{F}[f(t,.)]-\mathcal{F}_{*}\geq 0$ and  $\mathcal{F}[f]$ is nonincreasing, this proves that
$\lim_{t\rightarrow\infty}\mathcal{F}[f(t,.)]=\mathcal{F}_{*}.$
Using (13) and Lemma˜3.2, we obtain
$0\leq\lim_{t\rightarrow\infty}D\,\mathcal{H}[f(t,.)|G_{u_{f}(t)}]\leq\lim_{t\rightarrow\infty}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})=0.$
Hence, there exists a time $t_{1}>t_{0}$ such that
$\varphi\left(\mathcal{H}[f(t,.)|G_{u_{f}(t)}]\right)\leq(1+2^{\frac{1}{p}}A_{q})\mathcal{H}[f(t,.)|G_{u_{f}(t)}]^{\frac{1}{p}}\,\,\,\forall t\geq t_{1}.$ (27)
Using (26), (25) and (27), we obtain
$\frac{d}{dt}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})\leq C(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})^{p}\,\,\,\,\forall t\geq t_{1},$ (28)
for some constant $C>0$. Integrating (28) with respect to $t\geq t_{1}$, we obtain the result with $\beta=1/(p-1)$.
3.3. Case $D<D_{*}$
[…]
@@ 3.3. Case $D<D_{*}$
Let $D\geq D_{*}$ and let $f$ be the solution to the Equation˜1 with initial datum  $f_{\mathrm{ini}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{ini}}]<+\infty$ and $\|f_{\mathrm{ini}}\|_{L_{1}(\mathbb{R}^{d})}=1$. For all $\beta>0$ if $D>D_{*}$ and all $\beta\in(0,2)$ if $D=D_{*}$, then
@@
    A log-Sobolev inequality holds for $G_{u}$ where $u$ belongs to a neighborhood of 0 and it is also the case since $\log(G_{0})$ is a bounded perturbation of a uniformly convex potential.
@@
    $u_{f}$ converges to 0 according to the case $D\geq D_{*}$ of Lemma˜2.2.
@@
Doing the same computations as in Proposition˜3.1, we conclude with $\beta=1/(p-1)$ for $D>D_{*}$ and with $\beta=2/(3p-2)$ for $D=D_{*}$. ∎
@@
By Lemma˜2.2, we know that $\lim_{t\rightarrow\infty}u_{f}(t)=0$. Using a Taylor expansion near 0, using Lemma˜3.4, we obtain that there exist positive constants $C_{1}$ and $C_{2}$ such that for $t$ large enough,
$|u_{f}|^{2}\leq C_{1}\left(\mathcal{V}(u_{f})-\mathcal{V}(0)\right)\leq C_{1}\left(\mathcal{F}(f)-\mathcal{F}(G_{0})\right)\text{ if }D>D_{*},$
$|u_{f}|^{4}\leq C_{2}\left(\mathcal{V}(u_{f})-\mathcal{V}(0)\right)\leq C_{2}\left(\mathcal{F}(f)-\mathcal{F}(G_{0})\right)\text{ if }D=D_{*},$
where we used the inequality $\mathcal{V}(u_{f})-\mathcal{V}(0)\leq\mathcal{F}(f)-\mathcal{F}(G_{0})$ which is a consequence of (13). This concludes the proof using Proposition˜3.2 and (15). ∎
4. Convergence of $u_{f}$ and proof of Theorem˜1.1
@@
 $\lim_{t\rightarrow\infty}{\mathrm{dist}(u_{f},\mathcal{S})}=0$, we deduce from (15) that $\lim_{t\rightarrow\infty}\inf_{u\in\mathcal{S}}\mathcal{H}[f(t,.)|G_{u}]=0$. We now want to prove that $f$ converges to a unique stationary solution (in relative entropy). Using (15) this is equivalent to proving that $u_{f}$ converges. We proved in Section˜2.7 that $u_{f}$ converges to $\mathcal{S}$, which means that $\lim_{t\rightarrow\infty}|u_{f}(t)|=r(D),$ but we did not prove yet that $u_{f}(t)$ converges as $t\rightarrow+\infty$ to some fixed $u\in\mathcal{S}$.
@@
Under the assumptions of Theorem˜1.1, there exists $u_{\infty}\in\mathcal{S}$ such that for all $\beta>0$,
$\lim_{t\rightarrow\infty}t^{\beta}|u_{f}(t)-u_{\infty}|=0.$
@@
$\frac{du_{f}}{dt}=\alpha u_{f}-\alpha\int_{\mathbb{R}^{d}}v|v|^{2}fdv$ (29)
and that for all $u\in\mathcal{S}$ we have $u=\int_{\mathbb{R}^{d}}|v|^{2}\,vG_{u}\,dv$, we obtain
$\left|\frac{du_{f}}{dt}(t)\right|$ $\leq\alpha\left|u_{f}-u\right|+\alpha\int_{\mathbb{R}^{d}}|v|^{3}\,|f-G_{u}|\,dv$
$\leq\alpha\left|u_{f}-u\right|+\alpha\left(\int_{\mathbb{R}^{d}}\,|v|^{3q}|f-G_{u}|\,dv\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{d}}|f-G_{u}|dv\,\right)^{\frac{1}{p}}$
$\leq\alpha\left|u_{f}(t)-u\right|+C_{q}\,\lVert f(t,.)-G_{u}\rVert_{L^{1}(\mathbb{R}^{d})}^{\frac{1}{p}}.$
@@
$C_{q}=\sup_{t>t_{0}}{\alpha\left(\int_{\mathbb{R}^{d}}|v|^{3q}|f-G_{u}|dv\right)^{\frac{1}{q}}}<+\infty$
because $f$ has finite moments for $t>0$ by Lemma˜2.1. Since this is true for all $u\in\mathcal{S}$, we can choose
$u_{*}(t)=r(D)\frac{u_{f}(t)}{|u_{f}(t)|}.$ (30)
Since $\lim_{t\rightarrow\infty}\mathrm{dist}(u_{f},\mathcal{S})=0$, there exists $t_{0}>0$ such that for all $t\geq t_{0}$, $u_{f}(t)\in\mathcal{N}$ where $\mathcal{N}$ is a neighborhood of $\mathcal{S}$ where (19) applies. Hence, $\forall t\geq t_{0}$,
$|u_{f}(t)-u_{*}(t)|^{2}=\mathrm{dist}(u_{f}(t),\mathcal{S})^{2}\leq\frac{2}{\mu_{1}}(\mathcal{V}(u_{f}(t))-\mathcal{V}_{*})\leq\frac{2}{\mu_{1}}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})$
@@
$||f(t,.)-G_{u_{*}(t)}||_{L^{1}(\mathbb{R}^{d})}^{\frac{1}{p}}\leq\mathcal{H}[f(t,.)|G_{u_{*}(t)}]^{\frac{1}{2p}}\leq\left(\frac{1}{D}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})+\frac{1}{2D}|u_{f}(t)-u_{*}(t)|^{2}\right)^{\frac{1}{2p}}$

### v3

@@ 2.6. Log-Sobolev and Poincaré inequality
$\frac{d}{dt}\mathcal{F}[f(t,.)]=-\mathcal{I}[f(t,.)]\leq-\,D^{2}\,\mathcal{K}\,\mathcal{H}[f|G_{\mathbf{u}_{f}}],$
where $\mathcal{K}=\mathcal{K}_{\epsilon}$ is the positive constant in (LSI) with $\epsilon=\sup_{t\geq 0}\mathrm{dist}(\mathbf{u}_{f}(t),\mathcal{S})<+\infty$ by Lemma˜2.1. Since $t\mapsto\mathcal{F}[f(t,.)]$ is nonincreasing and bounded from below, it has a limit $\mathcal{F}_{\infty}\in\mathbb{R}$ as $t\rightarrow+\infty$. Integrating the previous inequality in time, we get a de Bruijn type estimate
$\int_{0}^{\infty}\mathcal{H}[f|G_{\mathbf{u}_{f}}]dt\leq\frac{1}{\mathcal{K}D^{2}}(\mathcal{F}[f_{\mathrm{in}}]-\mathcal{F}_{\infty}).$ (17)
With (16), (C-K) and (17), we obtain that $\gamma(t)=W(\mathbf{u}_{f}(t))$ is integrable on $\mathbb{R}_{+}$, where $W$ is defined on $\mathbb{R}^{d}$ by $W(Y)=|Y-\mathbf{u}_{G_{Y}}|^{\beta}$. Observe that $\gamma^{\prime}(t)=\nabla W(\mathbf{u}_{f}(t))\mathbf{u}_{f}^{\prime}(t)$ is bounded since $\mathbf{u}_{f}$ and $\mathbf{u}_{f}^{\prime}=\alpha\mathbf{u}_{f}-\alpha\int_{\mathbb{R}^{d}}v|v|^{2}f\,dv$ are uniformly bounded in $t\geq 0$. Since $\gamma$ is integrable in time and uniformly continuous, we obtain $\lim_{t\rightarrow\infty}\gamma(t)=0$ thanks to Barbalat’s lemma. Using (8), we obtain :
$|\mathbf{u}_{f}(t)-\mathbf{u}_{G_{\mathbf{u}_{f}}}(t)|=-\,\frac{{H}(|\mathbf{u}_{f}|)}{{K}(|\mathbf{u}_{f}|)}\frac{\mathbf{u}_{f}}{|\mathbf{u}_{f}|}.$
@@
$\lim_{t\rightarrow{\infty}}\mathbf{u}_{f}(t)=0\,.$
If $D<D_{*}$, we recall that the zeros of H{H} on $\mathbb{R}_{+}$ are exactly $r(D)$ and 0 by Proposition˜1.1. If $\mathcal{F}[f_{\mathrm{in}}]<\mathcal{F}[G_{0}]$, the case $\lim_{t\rightarrow{\infty}}\mathbf{u}_{f}(t)=0$ is excluded, since in this case
$\mathcal{F}[f_{\mathrm{in}}]\geq\lim_{t\rightarrow\infty}\mathcal{F}[f]\geq\inf_{\mathbf{u}_{f}=0}\mathcal{F}[f]=\mathcal{F}[G_{0}]>\mathcal{F}[f_{\mathrm{in}}].$
@@
$\lim_{t\rightarrow{\infty}}\mathrm{dist}(\mathbf{u}_{f},\mathcal{S})=0\,.$
@@
$\mathcal{V}(\mathbf{u})=-\,D\log\left(\int_{\mathbb{R}^{d}}e^{-\frac{1}{D}\left(\frac{1}{2}|v-\mathbf{u}|^{2}+\psi_{\alpha}(v)\right)}\,dv\right).$ (18)
In particular, $\mathcal{V}$ is smooth and radial and with a slight abuse of notation, we shall write $\mathcal{V}(\mathbf{u})=\mathcal{V}(|\mathbf{u}|)$. The key point is to prove estimates of Polyak–Łojasiewicz type, which are based on the following Taylor expansion.
@@
  • If $D<D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{\prime\prime}(r(D))|<\epsilon$ for $i=1,2$ and $0<\delta<\mathrm{min}\big\{r(D)/2,\epsilon\big\}$ such that for all $\mathbf{u}$ satisfying $\mathrm{dist}(\mathbf{u},\mathcal{S})<\delta$ we have,
    $\frac{\mu_{1}}{2}\mathrm{dist}(\mathbf{u},\mathcal{S})^{2}\leq\mathcal{V}(\mathbf{u})-\mathcal{V}_{*}\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(\mathbf{u})|^{2}.$ (19)
  • If $D>D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{\prime\prime}(0)|<\epsilon$ for $i=1,2$ and $0<\delta<\epsilon$ such that for all $|\mathbf{u}|<\delta$ we have,
    $\frac{\mu_{1}}{2}|\mathbf{u}|^{2}\leq\mathcal{V}(\mathbf{u})-\mathcal{V}(0)\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(\mathbf{u})|^{2}.$ (20)
  • If $D=D_{*}$. For all $\epsilon>0$, there exist $\mu_{1}>0$ and $\mu_{2}>0$ which satisfy $|\mu_{i}-\mathcal{V}^{(4)}(0)|<\epsilon$ for $i=1,2$ and $0<\delta<\epsilon$ such that for all $|\mathbf{u}|<\delta$ we have,
    $\frac{\mu_{1}}{24}|\mathbf{u}|^{4}\leq\mathcal{V}(\mathbf{u})-\mathcal{V}(0)\leq\ \frac{6^{\frac{4}{3}}}{24\mu_{2}^{\frac{1}{3}}}|\nabla\mathcal{V}(\mathbf{u})|^{\frac{4}{3}}.$ (21)
@@
The proof follows from a Taylor expansion of $\mathcal{V}$ and $\mathcal{V}^{\prime}$ around $\mathbf{u}\in\mathcal{S}$ if $D<D_{*}$ and around 0 if $D\geq D_{*}$. Since $\mathcal{S}$ and 0 are critical points of $\mathcal{V}$ (see Lemma˜3.2), one has to check that:
@@
It is shown in the proof of [11, Proposition 2.3] for the case $d\geq 2$, and in the proof of [11, Proposition 2.2] for $d=1$, that
@@
We will prove that $H^{\prime}$ has a unique zero $\mathbf{u}_{0}$ and $0<\mathbf{u}_{0}<r(D)$. Since $D<D_{*}$, thanks to Proposition˜1.1, we know that $H(0)=0$ and $H(r(D))=0$ with $r(D)>0$. Using Rolle’s lemma, we conclude that $H^{\prime}$ has at least one zero on $(0,r(D))$.
@@ 3.2. Case $D<D_{*}$
Let $\epsilon>0$. By Lemma˜3.1 there exist positive constants $\mu_{1}$ and $\mu_{2}$ such that $|\mu_{i}-\mathcal{V}^{\prime\prime}(r(D))|<\epsilon$ and a neighborhood $\mathcal{N}=\mathcal{N}_{\epsilon}$ of $\mathcal{S}$ such that (19) holds for all $\mathbf{u}\in\mathcal{N}$. By Lemma˜2.2, that is $\lim_{t\rightarrow\infty}\mathrm{dist}(\mathbf{u}_{f}(t),\mathcal{S})=0$, there exists $t_{0}>0$ such that for all $t\geq t_{0}$ we have $\mathbf{u}_{f}(t)\in\mathcal{N}$. Hence, we have,
$\mathcal{V}(\mathbf{u}_{f}(t))-\mathcal{V}_{*}\leq\frac{1}{2\mu_{2}}|\nabla\mathcal{V}(\mathbf{u}_{f}(t))|^{2}$ $=\frac{1}{2\mu_{2}}|\mathbf{u}_{f}-\mathbf{u}_{G_{\mathbf{u}_{f}}}|^{2}$
@@
$\leq A_{q}(2\mathcal{H}[f(t,.)|G_{\mathbf{u}_{f}(t)}])^{\frac{1}{p}}\text{ for all }t\geq t_{0},$ (25)
@@
$A_{q}=\frac{1}{2\mu_{2}}\sup_{t>t_{0}}{\left(\int_{\mathbb{R}^{d}}|v|^{q}|f+G_{\mathbf{u}_{f}}|\,dv\right)^{\frac{2}{q}}}$
@@
$\mathcal{F}[f]-\mathcal{F}_{*}\leq\varphi(\mathcal{H}[f|G_{\mathbf{u}_{f}}])\,\,\,\forall t\geq t_{0}$ (26)
@@
$\frac{d}{dt}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})=-\mathcal{I}[f(t,.)]\leq-D^{2}\mathcal{K}\mathcal{H}[f(t,.)|G_{\mathbf{u}_{f}(t)}]\leq-D^{2}\mathcal{K}\varphi^{-1}\left(\mathcal{F}[f(t,.)]-\mathcal{F}_{*}\right),$ (27)
using (LSI) with constant $\mathcal{K}=\mathcal{K}_{\epsilon}$ since $\mathrm{dist}(\mathbf{u}_{f}(t),\mathcal{S})<\epsilon$ for all $t\geq t_{0}$. Since $\mathcal{F}[f(t,.)]-\mathcal{F}_{*}\geq 0$ and $\mathcal{F}[f]$ is nonincreasing, this proves that
$\lim_{t\rightarrow\infty}\mathcal{F}[f(t,.)]=\mathcal{F}_{*}.$ (28)
Using (27) and (28), we obtain that there exists $t_{1}\geq t_{0}$
$\frac{d}{dt}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})\leq-C(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})^{p}\,\,\,\,\forall t\geq t_{1},$ (29)
for some constant $C>0$. Integrating (29) with respect to $t\geq t_{1}$, we obtain the result with $\beta<1/(p-1)$.
3.3. Case $D\geq D_{*}$
In [11], the exponential decay of $f$ to $G_{0}$ with respect to the $L^{2}(\mathbb{R}^{d},G_{0}^{-1}dv)$ is proved for $D>D_{*}$ and for initial data $f_{\mathrm{in}}\in L^{2}(\mathbb{R}^{d},G_{0}^{-1}dv)$. In this section, following the same method as in Section˜3.2, we obtain that the convergence of $f$ to $G_{0}$ (in relative entropy) is still true by only assuming $f_{\mathrm{in}}\in L^{1}(\mathbb{R}^{d})$ and $\mathcal{F}[f_{\mathrm{in}}]<+\infty$. We only get an arbitrary algebraic rate of convergence. The limit case $D=D_{*}$ is also of interest. We know that for $D=D_{*}$, Equation˜1 has a unique stationary state $G_{0}$. To our knowledge, there is no convergence result in the literature even under the condition $f_{\mathrm{in}}\in L^{2}(G_{0}^{-1})$. We are going to prove this convergence in relative entropy.
@@ 3.3. Case $D\geq D_{*}$
Let $D\geq D_{*}$ and let $f$ be the solution to the Equation˜1 with initial datum $f_{\mathrm{in}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{ini}}]<+\infty$ and $\|f_{\mathrm{in}}\|_{L_{1}(\mathbb{R}^{d})}=1$. For all $\beta>0$ if $D>D_{*}$ and all $\beta\in(0,2)$ if $D=D_{*}$, then
@@
    A log-Sobolev inequality holds for $G_{\mathbf{u}}$ where $\mathbf{u}$ belongs to a neighborhood of 0 and it is also the case since $\log(G_{0})$ is a bounded perturbation of a uniformly convex potential.
@@
    $\mathbf{u}_{f}$ converges to 0 according to the case $D\geq D_{*}$ of Lemma˜2.2.
@@
Doing the same computations as in Proposition˜3.1, we conclude with $\beta<1/(p-1)$ for $D>D_{*}$ and with $\beta<2/(3p-2)$ for $D=D_{*}$. ∎
@@
By Lemma˜2.2, we know that $\lim_{t\rightarrow\infty}\mathbf{u}_{f}(t)=0$. Using a Taylor expansion near 0, using Lemma˜3.4, we obtain that there exist positive constants $C_{1}$ and $C_{2}$ such that for $t$ large enough,
$|\mathbf{u}_{f}|^{2}\leq C_{1}\left(\mathcal{V}(\mathbf{u}_{f})-\mathcal{V}(0)\right)\leq C_{1}\left(\mathcal{F}(f)-\mathcal{F}(G_{0})\right)\text{ if }D>D_{*},$
$|\mathbf{u}_{f}|^{4}\leq C_{2}\left(\mathcal{V}(\mathbf{u}_{f})-\mathcal{V}(0)\right)\leq C_{2}\left(\mathcal{F}(f)-\mathcal{F}(G_{0})\right)\text{ if }D=D_{*},$
where we used the inequality $\mathcal{V}(\mathbf{u}_{f})-\mathcal{V}(0)\leq\mathcal{F}(f)-\mathcal{F}(G_{0})$ which is a consequence of (13). This concludes the proof using Proposition˜3.2 and (15). ∎
4. Convergence of $\mathbf{u}_{f}$ and proof of Theorem˜1.1
@@
$\lim_{t\rightarrow\infty}{\mathrm{dist}(\mathbf{u}_{f},\mathcal{S})}=0$, we deduce from (15) that $\lim_{t\rightarrow\infty}\inf_{\mathbf{u}\in\mathcal{S}}\mathcal{H}[f(t,.)|G_{\mathbf{u}}]=0$. We now want to prove that $f$ converges to a unique stationary solution (in relative entropy). Using (15) this is equivalent to proving that $\mathbf{u}_{f}$ converges. We proved in Section˜2.7 that $\mathbf{u}_{f}$ converges to $\mathcal{S}$, which means that $\lim_{t\rightarrow\infty}|\mathbf{u}_{f}(t)|=r(D),$ but we did not prove yet that $\mathbf{u}_{f}(t)$ converges as $t\rightarrow+\infty$ to some fixed $\mathbf{u}\in\mathcal{S}$.
@@
Under the assumptions of Theorem˜1.1, there exists $\mathbf{u}_{\infty}\in\mathcal{S}$ such that for all $\beta>0$,
$\lim_{t\rightarrow\infty}t^{\beta}|\mathbf{u}_{f}(t)-\mathbf{u}_{\infty}|=0.$
@@
$\frac{d\mathbf{u}_{f}}{dt}=\alpha\mathbf{u}_{f}-\alpha\int_{\mathbb{R}^{d}}v|v|^{2}fdv$ (30)
and that for all $\mathbf{u}\in\mathcal{S}$ we have $\mathbf{u}=\int_{\mathbb{R}^{d}}|v|^{2}\,vG_{\mathbf{u}}\,dv$, we obtain
$\left|\frac{d\mathbf{u}_{f}}{dt}(t)\right|$ $\leq\alpha\left|\mathbf{u}_{f}-\mathbf{u}\right|+\alpha\int_{\mathbb{R}^{d}}|v|^{3}\,|f-G_{\mathbf{u}}|\,dv$
$\leq\alpha\left|\mathbf{u}_{f}-\mathbf{u}\right|+\alpha\left(\int_{\mathbb{R}^{d}}\,|v|^{3q}|f-G_{\mathbf{u}}|\,dv\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{d}}|f-G_{\mathbf{u}}|dv\,\right)^{\frac{1}{p}}$
$\leq\alpha\left|\mathbf{u}_{f}(t)-\mathbf{u}\right|+C_{q}\,\lVert f(t,.)-G_{\mathbf{u}}\rVert_{L^{1}(\mathbb{R}^{d})}^{\frac{1}{p}}.$
@@
$C_{q}=\sup_{t>t_{0}}{\alpha\left(\int_{\mathbb{R}^{d}}|v|^{3q}|f-G_{\mathbf{u}}|dv\right)^{\frac{1}{q}}}<+\infty$
because $f$ has finite moments for $t>0$ by Lemma˜2.1. Since this is true for all $\mathbf{u}\in\mathcal{S}$, we can choose
$\mathbf{u}_{*}(t)=r(D)\frac{\mathbf{u}_{f}(t)}{|\mathbf{u}_{f}(t)|}.$ (31)
Since $\lim_{t\rightarrow\infty}\mathrm{dist}(\mathbf{u}_{f},\mathcal{S})=0$, there exists $t_{0}>0$ such that for all $t\geq t_{0}$, $\mathbf{u}_{f}(t)\in\mathcal{N}$ where $\mathcal{N}$ is a neighborhood of $\mathcal{S}$ where (19) applies. Hence, $\forall t\geq t_{0}$,
$|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|^{2}=\mathrm{dist}(\mathbf{u}_{f}(t),\mathcal{S})^{2}\leq\frac{2}{\mu_{1}}(\mathcal{V}(\mathbf{u}_{f}(t))-\mathcal{V}_{*})\leq\frac{2}{\mu_{1}}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})$
@@
$||f(t,.)-G_{\mathbf{u}_{*}(t)}||_{L^{1}(\mathbb{R}^{d})}^{\frac{1}{p}}\leq\mathcal{H}[f(t,.)|G_{\mathbf{u}_{*}(t)}]^{\frac{1}{2p}}\leq\left(\frac{1}{D}(\mathcal{F}[f(t,.)]-\mathcal{F}_{*})+\frac{1}{2D}|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|^{2}\right)^{\frac{1}{2p}}$